OpenStudy (anonymous):

A parachutist's speed during a free fall reaches 226 kilometers per hour. What is this speed in meters per second? At this speed, how many meters will the parachutist fall during 15 seconds of free fall? Do not round your answers

OpenStudy (tkhunny):

\(\dfrac{226\;km}{hr}\cdot\dfrac{1\;hr}{60\;min}\dfrac{1\;min}{60\;sec}\) There's km/sec. Now what?

OpenStudy (anonymous):

i dont know

OpenStudy (tkhunny):

Convert from km to m. Show us your metric chops!!

OpenStudy (anonymous):

226000 meters

OpenStudy (tkhunny):

That's now meters per hour. Go add that conversion factor to the two I provided: \(\dfrac{1000\;m}{1\;km}\)

OpenStudy (anonymous):

am i suppose to multiply the three you provide above?

OpenStudy (tkhunny):

Once you get all the conversion factors lines up nicely, yes. Watch how all the undesirable units cancel each other out and freely multiply numerators and denominators.